find the length and width of a rectangular lot with a perimeter of 88 meters if the length is 4 meters more than the width.
Answer by checkley77(12844)

p=2l*2w
88=2(w+4)+2w
88=2w+8+2w
88=4w-8
4w=88-8
4w=80
w=80/4
w=20 meters for the width.
l=20+4=24 meters for the length.
proof:
2*24+2*20=88
48+40=88
88=88